# Joint Information Extraction Across Classical and Modern Chinese with Tea-MOELoRA

### v1

@@ Abstract
Xuemei Tang1, Chengxi Yan2thanks: Corresponding authors., Jinghang Gu1thanks: Corresponding authors., and Chu-Ren Huang1 1The Hong Kong Polytechnic University, Hong Kong, China 2Renmin University of China, Beijing, China xuemeitang00@gmail.com
@@ 4.5 Effect of Router Signal Merge Strategy
4.6 Expert Utilization Preferences of Tasks and Eras
We present the visualization of expert weight distribution in Figure LABEL:weight_districution. According to Figure LABEL:f3a, the results demonstrate that samples from different eras are indeed assigned to different experts. Specifically, Classical Chinese samples are predominantly assigned to Expert 4 and Expert 7, whereas Modern Chinese samples are mainly assigned to Expert 0 and Expert 1. For each task, Figure LABEL:f3b shows that different tasks emphasize different experts. Specifically, CHisRE mainly relies on Expert 5, DUIE primarily attends to Expert 0 and Expert 1, CHED is focused on Expert 7, and ACE is assigned mainly to Expert 2. These results indicate that both era and task factors guide the model to make distinct expert selections, confirming the effectiveness of the expert specialization mechanism.

### v2

@@ Abstract
Xuemei Tang1, Chengxi Yan2thanks: Corresponding authors., Jinghang Gu1111, and Chu-Ren Huang1 1The Hong Kong Polytechnic University, Hong Kong, China 2Renmin University of China, Beijing, China xuemeitang00@gmail.com
@@ 4.5 Effect of Router Signal Merge Strategy
4.6 Expert Utilization Preferences
Figure 6: Case study of relation extraction and event detection. Green highlights indicate hallucinated predictions, while blue highlights denote incorrect predictions.
We present the visualization of expert weight distribution in Figure LABEL:weight_districution. According to Figure LABEL:f3a, the results demonstrate that samples from different eras are indeed assigned to different experts. Specifically, Classical Chinese samples are predominantly assigned to Expert 4 and Expert 7, whereas Modern Chinese samples are mainly assigned to Expert 0 and Expert 1. For each task, Figure LABEL:f3d shows that different tasks emphasize different experts. Specifically, CHisRE mainly relies on Expert 5, DUIE primarily attends to Expert 0 and Expert 1, CHED is focused on Expert 7, and ACE is assigned mainly to Expert 2. These results indicate that both era and task factors guide the model to make distinct expert selections, confirming the effectiveness of the expert specialization mechanism.
Moreover, by comparing Figures LABEL:f3aLABEL:f3b,  LABEL:f3cLABEL:f3d and the smoothness metrics (variance, entropy, and max-min difference as shown in Appendix A), we observe that the distribution for models using only Era ID or Task ID is generally smoother and more balanced across experts, while the combined Era and Task ID model shows more concentrated assignments, indicating stronger specialization when both factors are considered.
4.7 Effect of Key Parameter
4.8 Case Study
We present two representative examples to illustrate the performance differences among Tea-MoELORA(Ours), MoELORA, TeamLoRA, and LoRA (Mix) on relation extraction and event detection.
In the first example, from CHisRE, the sentence mentions the officials “胡土门”(Hutumen) and “毛花辇” (Maohuanian), and their respective positions. Our model correctly predicts all three relations: the official positions of “胡土门” and “毛花辇”, and their colleague relationship. Other models either hallucinate additional relations, such as management relations involving places “河朔 (Heshuo)”or incorrectly link “胡土门 (Hutumen)” to other entities, demonstrating issues with over-prediction or missing entity constraints.
In the second example, from ACE, the sentence describes a conflict between Palestinian and Israeli parties. Our model accurately extracts the trigger “冲突” (Conflict), identifies the event type as “冲突-攻击” (Conflict-Attack), and assigns correct roles to the participants. In contrast, MoELORA, TeamLoRA, and LoRA (Mix) incorrectly merge the participants as a single entity (“巴以双方 (both Palestinian and Israeli)”), resulting in inaccurate argument assignments.
These examples highlight that our model not only reduces hallucinated relations but also better preserves entity-level accuracy and role assignments, especially in complex historical or conflict-related sentences.
@@ 5 Conclusion
Appendix A Appendix A
Model Variance Entropy Max-Min Diff
Task ID (Tea-MoELoRA) 0.01825 1.62666 0.40703
Era ID (Tea-MoELoRA) 0.00981 1.78265 0.23610
Task ID (only) 0.00557 1.90709 0.21752
Era ID (only) 0.00286 1.97587 0.16970
Table 5: Comparison of Heatmap Smoothness Across Different Models. Higher entropy and lower variance/Max-Min indicate a smoother distribution.